you are cooking rice for p people.
each bag of rice serves three people.
and you need four extra bags for safe measure.
if you have 8 bags of rice what is the maximum amount of people you could serve for.

Answer by Theo(13342)

if each bag of rice serves 3 people, then each person is served by 1/3 of a bag of rice.

1/3 * p + 4 = 8

subtract 4 from both sides of this equation to get 1/3 * p = 4

multiply both sides of this equation by 3 to get p = 12.

you can serve 12 people with 8 bags of rice.

take 4 away and keep them for safe measure.

you will be using 4 bags of rice.

since each bag of rice can serve 3 people, 4 * 3 = 12 people that can be served by those 4 bags of rice.